0. The table below shows data for 13 students in a statistics class. Each member of the class ran a 40-yard sprin and then did a long iump (with a running start). Sprint Time (s) 5.41 5.05 9.49 8.09 7.01 7.17 6.83 6.73 8.01 5.68 5.78 6.31 6.04 Long Jump (in) 171 184 48 151 90 65 94 78 71 130 173 143 141 a) Create and label a scatterplot of the data (Sprint time vs Long Jump) b) Describe and interpret the scatterplot above. c) Calculate the LSRL equation and the regression coefficient r and R2. What can you say about them? d) Plot the Residuals vs Long jump. e) Calculate the LSRL and regression coefficient of the Residuals.
Answer #1

a)

b)

Scatterplot shows a negative direction

AS the values of sprint time,X increases, long jump distance,Y gets decrease

c)

X Y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
5.41 171 1.7648101 2768.379 -69.8975
5.05 184 2.8509024 4305.379 -110.789
9.49 48 7.5709639 4953.994 -193.666
8.09 151 1.8266562 1063.763 44.08095
7.01 90 0.0737331 805.6864 -7.70751
7.17 65 0.1862254 2849.917 -23.0375
6.83 94 0.0083793 594.6095 -2.23213
6.73 78 7.16E-05 1630.917 0.341716
8.01 71 1.6168101 2245.302 -60.2514
5.68 130 1.1203408 134.9172 -12.2944
5.78 173 0.9186485 2982.84 -52.3467
6.31 143 0.1835793 605.9172 -10.5467
6.04 141 0.4878485 511.4556 -15.796

.

ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 87.6 1539 18.60897 25453.08 -514.142
mean 6.738461538 118.384615 SSxx SSyy SSxy

sample size ,   n =   13          
here, x̅ =   6.738461538       ȳ =   118.3846154  
                  
SSxx =    Σ(x-x̅)² =    18.60896923          
SSxy=   Σ(x-x̅)(y-ȳ) =   -514.1423077          
                  
slope ,    ß1 = SSxy/SSxx =   -27.62873651          
                  
intercept,   ß0 = y̅-ß1* x̄ =   304.5597937          
                  
so, regression line is   Ŷ =   304.5598   +   -27.6287   *x
                  
SSE=   (Sx*Sy - S²xy)/Sx =    11247.97          
                  
std error ,Se =    √(SSE/(n-2)) =    31.9772          
                  
correlation coefficient ,    r = Sxy/√(Sx.Sy) =   -0.7471          
                  
R² =    (Sxy)²/(Sx.Sy) =    0.5581          
                  
there is negative correlation between X and Y

and 55.81% variation in variable Y is explained by variable X

d)

Y Residual
171 155.0883 15.91167
184 165.0347 18.96533
48 42.36308 5.636916
151 81.04332 69.95668
90 110.8824 -20.8824
65 106.4618 -41.4618
94 115.8555 -21.8555
78 118.6184 -40.6184
71 83.25361 -12.2536
130 147.6286 -17.6286
173 144.8657 28.1343
143 130.2225 12.77753
141 137.6822 3.317775

please note that only first four subparts of a ques per post can be answered as per HOMEWORKLIB RULES